OpenStudy (anonymous):

SOMEONE PLEASE HELP ME ON MATH PROBLEMS PLEAASE IM IN ALG 2

OpenStudy (anonymous):

wut qusion

OpenStudy (anonymous):

Post them and we will try to help you.

OpenStudy (anonymous):

find the annual percent increase or decrease that y=0.35(2.3)x models

OpenStudy (anonymous):

0

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

0 persent decrese or incerese

OpenStudy (anonymous):

cool! i'm in algebra one but hey i'm only in 8th grade witch means i take online high school coarse that's pretty good right?

OpenStudy (anonymous):

im confused???

OpenStudy (anonymous):

i need help??

OpenStudy (anonymous):

i tell u da answr u no listen not mi folt

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

@Cheycheyy2 lets take this step by step okay :-). This is a exponential growth type question, have you seen these before or are you just learning them and are completely confused? Are you stuck anywhere specific?

OpenStudy (anonymous):

im horrible at math and im confused onthis whole section

OpenStudy (anonymous):

Okay, don't worry about if you don't understand your whole section, lets just focus on the question at hand. Exponential growth models this formula:\[\large x_t = x_0(1+r)^t\] Notice your question: \[\large y=0.35(2.3)^x\] Do you see how your question relates to the original form?

OpenStudy (anonymous):

yeah a little'

OpenStudy (anonymous):

Awesome :D. \[\large x_t = x_0(1+r)^t\] \[\large x_t \implies y\] \[\large x_0 \implies 0.35\] \[\large (1 + r) \implies 2.3 \] Notice the formula has a \((1+r)\). Your question doesn't, this is because your equation is written in final form. (Meaning the 1 and r (rate) are added together.) \[\large y=0.35(2.3)^x \implies \large y=0.35(2.3 - 1)^x \implies \large y=0.35(1.3)^x\] So \(1.3\) is our initial rate, do you know how to turn a decimal into a percent?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

its130%

OpenStudy (anonymous):

Correct! \[1.3 \times 100 = 130%\]

OpenStudy (anonymous):

thanks can you help me on more

OpenStudy (anonymous):

Sure, just post a new question :-). Anyway if you want to know more about exponential growth and decay, you can check out: http://www.regentsprep.org/regents/math/algebra/ae7/expdecayl.htm

OpenStudy (anonymous):

write expression as a single logarthim

OpenStudy (anonymous):